# Sample efficient likelihood-free inference for virus dynamics with different types of experiments

### v1

@@ 1 Introduction
We applied BOLFI to influenza A virus dynamics experimental data to learn the six parameters in the ordinary differential equation model. The target parameters are listed in table I. Previous research employed comprehensive MCMC-based approaches, requiring extensive computation time (e.g. 43 hours on a cluster to analyze 500,000 parameter sets)[11]. Moreover, this learning is based on small scale datasets. To improve the efficiency of learning, we applied the BOLFI algorithm to the same data sets and the same simulated data generating models with arrangement of combination. Remarkably, BOLFI completed this task with only 1 hour and 26 minutes for surrogate model learning, plus the sampling process, in total it is about 2 hours to obtain the posterior distributions for all six parameters, representing a 21.5-fold efficiency gain. Notably, our method accomplishes this without relying on explicit likelihood functions, thus frees researchers from extensive efforts on likelihood examinations previously necessary.
@@
  • Improvement of the BOLFI algorithm with Gaussian process based classifier for treatment of extreme values are provided.
@@ 2 Problem setting
Parameter Symbol
rate of one infectious virion entries a cell and successfully infects it $\gamma$
rate of infectious virion irreversible entry a cell $\beta$
infectious production rate for total RNA virions $P_{RNA}$
infectious production rate for infectious virions $P$
mean duration for the eclipse phase $\tau_{E}$
mean duration for the infectious phase $\tau_{I}$
Table I: Parameters to be inferred
@@ 2.2 The simulator
Employing the two simulators in [10], one describing the virus dynamics over time by a ordinary differential equation(ODE) and one simulating the ED assay experiment, we created a simulator which outputs four synthetic data sets(SC RNA data, SC ED data, MC RNA data and MC ED data) that have the same format as the experimental data that we are going to analyze. The ED assay simulator is stochastic, however, the ODE itself is deterministic. To make it a stochastic simulator, we added Gaussian random noise to the ODE simulated output and the variances of the noise are estimated from the experimental data to be learned. Adding noise to RNA measurements by scaling the RNA volume and applying a log-normal noise based on the specified standard deviation: $\sigma_{\text{MCRNA}}$ =0.226 characterizes variability in multiple-cycle total virus data; $\sigma_{\text{SCRNA}}$=0.145 characterizes variability in single-cycle total virus data. In [10], the ODE simulator and the ED assay simulator both only outputs one set of time series data. In contrast, we created a simulator repeats the simulation of one parameter set for $6$ times(because the largest repeats number in both the RNA and ED experimental data is $6$), then randomly select the same replica number for each data sets in each time step according to the experimental data from the $6$ sets of time series to create a simulated data has the same format of experimental data. Table IV in [10] summarizes all the fixed parameters and initial conditions used in simulating these infections. In appendix, A and B, we summarized the main explanation of the ODE simulator and ED assay simulator in [10]. Given the experimental data sets and the simulator for the four data sets, our goal is to infer the value of the parameters with uncertainty, $\gamma,\beta,p,p_{RNA},n_{I},n_{E}$ in the model.
@@ 3 The likelihood free inference method
3.1 The BOLFI algorithm[5]
@@
To mitigate the need for numerous potentially expensive simulations, active learning methods adapt the querying process based on different strategies. BOLFI employs Bayesian optimization [5] to iteratively build a probabilistic surrogate model of the relationship between parameters $\theta$ and discrepancies $d(\theta)=d(s(y),s(y_{\text{obs}}))$, where $s(.)$ is the summary statistics function over simulated data $y$ and observation data $y_{\text{obs}}$ and $d(.)$ represents a distance measure function, using the growing evidence set $\mathcal{E}_{t}$, which contains pairs $\{(\theta_{i},d(\theta_{i}))\}$ for $(i=1,\ldots,t)$.
@@
From $d(\theta)$, a suitable pointwise approximation to the likelihood function is retrieved as
@@
In practice, an acquisition function determines the locations of the relevant parameter space, and there are several reasonable choices. One example is the lower confidence bound selection criterion (LCBSC):
@@
where $\eta_{t}^{2}$ depends on the iteration $t$, the parameter space dimension $d$, and other tunable parameters [17]. The next parameter value to sample is obtained by minimizing (7) and randomly varying it to further balance the exploration and exploitation of the probabilistic function of the discrepancy.
3.2 BOLFI with Classifier(failure robust)–Treatment of extreme values
The simulator in [10] returns none output for invalid(physically not possible or not satisfying certain constrains) parameter combinations. In practice, the simulator uses several decision rules to check intermediate simulation results and decide whether to continue or abort the simulation. The rules can be viewed as hidden parameter constraints. While known parameter constraints such as $P_{RNA}>P$ by definition can be encoded in the prior distribution in BOLFI to avoid invalid parameter combinations and aborted simulations, this is not possible with the hidden constraints that are implemented in the simulator from [10] and not known in advance.
@@ 3.2 BOLFI with Classifier(failure robust)–Treatment of extreme values
The experiments carried out in this work use BOLFI extended with a classifier. The approach is similar to the one introduced by [4]. The idea is to combine the regression model that learns dependencies between simulator parameters and discrepancies with a binary classifier that learns the hidden constraints that divide parameters into valid and invalid combinations. While the regression model is only updated when the simulation succeeds and returns an output $y$, the classifier is updated based on all attempted and succeeded simulations, and predicted discrepancies are calculated based on the combined model. In practice, if the parameters $\theta$ are classified as valid, the predicted mean and variance are calculated based on Equations (3) and (4) as usual, and if the parameters are classified as invalid, the predicted mean $\mu(\theta)=\infty$ and predicted variance $\sigma(\theta)=0$.
@@
6:  if the simulator is valid for $\theta_{t}$ then
7:   GP classifier is classified as valid
8:   Simulate $y^{\prime}\sim p(y|\theta_{t})$
@@
12:   GP classifier is classified as invalid
@@
17:in the invalid region, set the predicted GP mean $\mu(\theta)=\infty$ and predicted variance $\sigma(\theta)=0$
18:
19:Learning:
20:for $t=N_{\text{init}}+1,\ldots,N_{E}$ do
21:  if $t\equiv T_{\text{update}}\mod T_{\text{update}}$ then
22:   Optimize GP hyperparameters
23:  end if
24:  Calculate $\theta_{t}=\arg\min A_{t}(\theta)$
25:  Generate $\theta_{t}\sim\mathcal{N}(\theta_{t},\sigma_{\text{acq}}^{2})$
26:  if the simulator is valid for $\theta_{t}$ then
27:   GP classifier is classified as valid
28:   Simulate $y^{\prime}\sim p(y|\theta_{t})$
29:   Calculate $d_{t}=d(s(y_{\text{obs}}),s(y^{\prime}))$
30:   Update $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup\{(\theta_{t},d_{t})\}$
31:   Fit $d(\theta)\mid\mathcal{E}_{t}\sim\text{GP}(\mu_{t}(\theta),v_{t}+\sigma^{2})$
32:  end if
33:  if the simulator is invalid for $\theta_{t}$ then
34:   GP classifier is classified as invalid
35:   in the invalid region, set the predicted GP mean $\mu(\theta)=\infty$ and predicted variance $\sigma(\theta)=0$
36:  end if
37:end for
38:
39:Output:
40:Calculate likelihood function $L(\theta)$ as in Equation (6)
41:Calculate corresponding posterior distribution with prior $\pi(\theta|y_{\text{obs}})=p(\theta)L(\theta)$
42:Draw samples from the learned posterior distribution $\pi(\theta|y_{\text{obs}})$
@@ 3.3 Setting
Adopting the same scale settings from previous studies, we set the parameter $\gamma$, $P$, beta($\beta$), $P_{RNA}$ to log scale as $\hat{\gamma}=\log_{10}{\gamma}=10^{\gamma}$, $\hat{P}=\log_{10}{P}=10^{P}$, $\hat{\beta}=\log_{10}{\beta}=10^{\beta}$, $\hat{P}_{RNA}=\log_{10}{P_{RNA}}=10^{P_{RNA}}$. We keep $\tau_{E}$ and $\tau_{I}$ on a linear scale.
@@
$P$ Uniform (log scale) $0.5$ to 33 Subject to $P<P_{RNA}$
$P_{RNA}$ Uniform (log scale) 22 to $4.8$ Subject to $P<P_{RNA}$
$\gamma$ Uniform (log scale) 1-1 to $0.5$
$\beta$ Uniform (log scale) 8-8 to 6-6
$\tau_{E}$ Uniform (linear scale) 22 to $15$
$\tau_{I}$ Uniform (linear scale) 22 to $55$
@@
Prior settings for parameters in the model are listed in table III. We added a condition for $P$ and $P_{RNA}$ as $P<P_{RNA}$. This condition is based on the definition of the parameters as $V$ is within $V_{RNA}$, see figure 2. First, we set prior for $P$ to be in log scale in bounds $(0.5,3)$. Prior for $P_{RNA}$ is in log scale in bounds $(2,4.8)$. Then we update parameter bounds adaptively during inference to create a dynamic prior distributions. We take the minimum value of $\hat{P}_{RNA}$ as the upper bound of $\hat{P}$ adaptively. This helps focus inference within a relevant parameter space, avoiding invalid regions to be more efficient. Figure 8 shows this constraint as “scale_p” which is the adaptive range of $\hat{P}$ and “upper_p” is the minimum value of $\hat{P}_{RNA}$.
@@ 4.2 Learning result from experimental data
Figure 10: This is a matrix of pairwise plots for Gaussian Process (GP) regression models(Contour plots use a blue color map), illustrating the relationships between parameters and their associated discrepancies. This figure visualize pairwise parameter relationships and their effects on discrepancies, also, provide a detailed visual representation of GP regression predictions and parameter dependencies. The red dots are samples. BOLFI+classifer inference with discrepancy defined in section 3.4 and Figure 8 and learning from experimental data.
@@ 5 Discussions and Conclusion
This study introduces an efficient likelihood-free inference framework for dynamical modeling of viral infection, leveraging the Bayesian Optimization for Likelihood-Free Inference (BOLFI) algorithm. By applying BOLFI to influenza A virus experimental data, we demonstrate that complex mechanistic models with high-dimensional parameter spaces can be inferred accurately and efficiently, even in the absence of explicitly used likelihood functions.
@@
Methodologically, we enhance the BOLFI framework with a GP-based classifier to handle extreme discrepancies and introduce a discrepancy design that integrates multiple types of experimental data. These innovations broaden the applicability of BOLFI to real-world biological inference problems involving heterogeneous data and partial model observability.
@@ Acknowledgments
This research was supported by the Finnish Center for Artificial Intelligence (FCAI). The project was initiated at iTHEMS RIKEN while Yingying Xu was supported by the SPDR program at RIKEN, Japan. Computational resources were provided by the CSC – IT Center for Science, Finland, through access to the Puhti supercomputing infrastructure. We thank Professor Jukka Corander and Professor Catherine Beauchemin for valuable discussions. We are also grateful to Professor Catherine Beauchemin and Dr. Christian Quirouette for explaining the problem and kindly providing the data and simulator from their earlier work.

### v2

@@ 1 Introduction
We aimed to infer the six parameters of an ordinary differential equation model describing influenza A virus dynamics, using experimental data. Traditionally, this problem has been addressed with comprehensive MCMC-based approaches, which require extensive computation time—often on the order of tens of hours (e.g., approximately 43 hours on a cluster for analyzing 500,000 parameter sets) [12]. Moreover, these methods are typically applied to small-scale datasets. To improve the efficiency of parameter inference, we applied the BOLFI algorithm to the same datasets and simulation models, with various arrangements. Remarkably, BOLFI completed the entire inference process—comprising surrogate model learning and sampling—in roughly 2 hours, using only 1,000 simulation runs, which is about a 20-fold reduction in computational effort (The actual times depend on the computing architecture; nevertheless, the computational time can provide an idea of the level of benefit achieved.). Notably, our method accomplishes this without relying on explicit likelihood functions, thus alleviating the extensive efforts previously required for likelihood examinations.
@@
  • Improvement of the BOLFI algorithm with Gaussian process based classifier for treatment of failed simulation cases are provided.
@@ 2 Problem setting
Symbol Parameter
$\gamma$ rate of one infectious virion entries a cell and successfully infects it
$\beta$ rate of infectious virion irreversible entry a cell
$P_{RNA}$ infectious production rate for total RNA virions
$P$ infectious production rate for infectious virions
$\tau_{E}$ mean duration for the eclipse phase
$\tau_{I}$ mean duration for the infectious phase
Table I: Parameters to be inferred.
@@ 2.2 The simulator
We developed a comprehensive simulator that models virus dynamics and experimental assay processes, based on two simulators described in [11]: one capturing the virus dynamics over time via an ordinary differential equation (ODE), and the other simulating the ED assay experiment. The simulator outputs four synthetic data sets(SC RNA data, SC ED data, MC RNA data, and MC ED data) that have the same format as the experimental data that we are going to analyze.
The ED assay simulator is stochastic, whereas the ODE itself is deterministic. To introduce stochasticity, Gaussian noise is added to the ODE output, with noise variances estimated from the experimental data:
$Y_{\text{RNA}}=V_{\text{RNA}}\times(1+\eta_{\text{RNA}}),\quad\eta_{\text{RNA}}\sim\mathcal{N}(0,\sigma^{2})$ (1)
where the noise variances are characterized as:
  • For multi-cycle RNA data: ($\sigma_{\text{MC RNA}}=0.226$)
  • For single-cycle RNA data: ($\sigma_{\text{SC RNA}}=0.145$)
To mimic repeated experimental measures, our combined simulator ($\mathcal{S}(\boldsymbol{\theta})$) repeats the simulation of each parameter set six times (matching the maximum number of replicates observed in the data). At each time point, it randomly selects one of these replicas according to the experimental replicating proportions:
$\mathcal{S}(\boldsymbol{\theta})$ $=\begin{aligned} &\left\{Y_{i}(t_{j})=\mathcal{R}_{I(t_{j},k)}^{(i)}(\boldsymbol{\theta})\;\middle|\;j=1,\dots,N_{t_{i}},\ k=1,\dots,M,\ i=\text{RNA, ED}\right\}\end{aligned}$ (2)
where $M$ is the maximum number of experimental replicates(in this case is set to $M=6$); $N_{t_{i}}$ is the maximum number of time steps in the corresponding experimental type i; $I(t_{j},k)$ represents the procedure of selecting a subset of $1,\dots,M$} which representing randomly selected index of replicates, reflecting the same replicated number with experimental data, independent for each $j$. This formulation ensures the synthetic data replicate the structure of real experimental datasets by incorporating variability both within and across replicates. In this way, we are generating synthetic datasets with the same structure as the experimental data. In contrast, in [11], the ODE simulator and the ED assay simulator both only output one set of time series data.
The combined simulator can be viewed as a single process with two interconnected compartments.
$\mathcal{S}(\boldsymbol{\theta})=\begin{cases}\text{Virus Dynamics (ODE)}\\ \text{Assay Simulation}\end{cases}$ (3)
where ($\boldsymbol{\theta}=(\gamma,\beta,P,P_{RNA},n_{I},n_{E}))$ represents the parameters to be inferred. In the Appendix (A and B), we provide detailed descriptions of the ODE-based virus dynamics model and the stochastic ED assay simulator. Table IV in [11] summarizes the fixed parameters and initial conditions used in the modeling. With this combined simulator and the experimental datasets, our goal is to infer the parameters ($\gamma,\beta,P,P_{RNA},n_{I},n_{E}$) and quantify their uncertainty.
@@ 3 The likelihood free inference method
3.1 The BOLFI algorithm (Gutmann and Corander, 2016)
@@
To mitigate the need for numerous potentially expensive simulations, active learning methods adapt the querying process based on different strategies. BOLFI employs Bayesian optimization [7] to iteratively build a probabilistic surrogate model of the relationship between parameters $\theta$ and discrepancies $d(\theta)=d(s(y),s(y_{\text{obs}}))$, where $s(.)$ is the summary statistics function over simulated data $y$ and observation data $y_{\text{obs}}$ and $d(.)$ represents a distance measurement function, using the growing evidence set $\mathcal{E}_{t}$, which contains pairs $\{(\theta_{i},d(\theta_{i}))\}$ for $(i=1,\ldots,t)$.
@@
From $d(\theta)$, a suitable pointwise approximation of the likelihood function is retrieved as
@@
In practice, an acquisition function determines the locations of the relevant parameter space, and there are several reasonable choices. The acquisition function used in this study is the maximum variance acquisition method [5]. In the maximum variance acquisition method, the next evaluation point is acquired in the maximizer of the variance of the unnormalized approximate posterior.
$\theta_{t+1}=\arg\max\text{Var}(p(\theta)\cdot p_{a}(\theta)),$ (10)
where $p(\theta)$ is the prior distribution, and the unnormalized likelihood $p_{a}$ is defined using the CDF for normal distribution, $\Phi(\cdot)$, as follows:
$p_{a}(\theta)=\Phi\left((\epsilon-\mu_{1:t}(\theta))/\sqrt{v_{1:t}(\theta)+\sigma^{2}_{n}}\right),$ (11)
where $\epsilon$ is the ABC threshold, $\mu_{1:t}$ and $v_{1:t}$ are determined by the Gaussian process, $\sigma^{2}_{n}$ is the noise variance which is set as $0.1$ in the inference. For more details refer to [5]. Other populor acquisition function, for example, the lower confidence bound selection criterion (LCBSC):
@@
where $\eta_{t}^{2}$ depends on the iteration $t$, the parameter space dimension $d$, and other tunable parameters [17]. The next parameter value to sample is obtained by minimizing (12) and randomly varying it to further balance the exploration and exploitation of the probabilistic function of the discrepancy. This acquisition function is not the one used in this study.
3.2 BOLFI with Classifier(failure robust)
The simulator in [11] fails with invalid(physically not possible or not satisfying certain constrains) parameter combinations. In practice, the simulator uses several decision rules to check intermediate simulation results and decide whether to continue or abort the simulation. The rules can be viewed as hidden parameter constraints. While known parameter constraints such as $P_{RNA}>P$ by definition can be encoded in the prior distribution in BOLFI to avoid invalid parameter combinations and aborted simulations, this is not possible with the hidden constraints that are implemented in the simulator from [11] and not known in advance. Especially, in complex systems, such as intermediate check constrains, those constrains are difficult to be included in prior settings. Therefore, we adopted the approach here in this study and believe it is useful for other situations as well when the simulator may fail in complex systems.
@@ 3.2 BOLFI with Classifier(failure robust)
The experiments carried out in this work use BOLFI extended with a classifier. The approach is similar to the one introduced by [6]. The idea is to combine the regression model that learns dependencies between simulator parameters and discrepancies with a binary classifier that learns the hidden constraints that divide parameters into valid and invalid combinations. While the regression model is only updated when the simulation succeeds and returns an output $y$, the classifier is updated based on all attempted and succeeded simulations, and predicted discrepancies are calculated based on the combined model. In practice, if the parameters $\theta$ are classified as valid, the predicted mean and variance are calculated based on Equations (6) and (7) as usual, and if the parameters are classified as invalid, the predicted mean $\mu(\theta)=\infty$ and predicted variance $\sigma(\theta)=0$. This setting of mean and variance in the prediction model is also helpful when selecting the next sampling point using the acquisition method, since it can avoid the invalid region.
@@
6:  Run simulator $y^{\prime}\sim p(y|\theta_{t})$
7:  if the simulator is valid for $\theta_{t}$ then
8:   Update GP classifier(classified as valid)
@@
12:   Update GP classifier(classified as invalid)
@@
17:
18:Learning:
19:for $t=N_{\text{init}}+1,\ldots,N_{E}$ do
20:  if $t\equiv T_{\text{update}}\mod T_{\text{update}}$ then
21:   Optimize regression GP hyperparameters
22:  end if
23:  Calculate $\theta_{t}$ based on acquisition method
24:  Run simulator $y^{\prime}\sim p(y|\theta_{t})$
25:  if the simulator is valid for $\theta_{t}$ then
26:   Update GP classifier(classified as valid)
27:   Calculate $d_{t}=d(s(y_{\text{obs}}),s(y^{\prime}))$
28:   Update $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup\{(\theta_{t},d_{t})\}$
29:   Fit $d(\theta)\mid\mathcal{E}_{t}\sim\text{GP}(\mu_{t}(\theta),v_{t}+\sigma^{2})$
30:  end if
31:  if the simulator is invalid for $\theta_{t}$ then
32:   Update GP classifier(classified as invalid)
33:  end if
34:end for
35:
36:Output:
37:Calculate likelihood function $L(\theta)$ as in Equation (9)
38:Calculate corresponding posterior distribution with prior $\pi(\theta|y_{\text{obs}})=p(\theta)L(\theta)$
39:Draw samples from the learned posterior distribution $\pi(\theta|y_{\text{obs}})$
@@ 3.3 Setting
Adopting the same scale settings from previous studies, we set the parameter $\gamma$, $P$, $\beta$, $P_{RNA}$ to log scale as $\hat{\gamma}=\log_{10}{\gamma}$, $\hat{P}=\log_{10}{P}$, $\hat{\beta}=\log_{10}{\beta}$, $\hat{P}_{RNA}=\log_{10}{P_{RNA}}$. We keep $\tau_{E}$ and $\tau_{I}$ on a linear scale.
@@
$\hat{P}$ Uniform (log scale) ($0.5$, 33) Subject to $\hat{P}<\hat{P}_{RNA}$
$\hat{P}_{RNA}$ Uniform (log scale) (22, $4.8$) Subject to $\hat{P}<\hat{P}_{RNA}$
$\hat{\gamma}$ Uniform (log scale) (1-1, $0.5$)
$\hat{\beta}$ Uniform (log scale) (8-8, 6-6)
$\tau_{E}$ Uniform (linear scale) (22, $15$)
$\tau_{I}$ Uniform (linear scale) (22, $55$)
@@
Prior settings for parameters in the model are listed in table III. We added a condition for $\hat{P}$ and $\hat{P}_{RNA}$ as $\hat{P}<\hat{P}_{RNA}$. This condition is based on the definition of the parameters as $V$ is within $V_{RNA}$, see figure 2. First, we set prior for $P$ to be in log scale in bounds $(0.5,3)$. Prior for $\hat{P}_{RNA}$ is in log scale in bounds $(2,4.8)$. Then we update parameter bounds adaptively during inference to create a dynamic prior distributions. We take the minimum value of $\hat{P}_{RNA}$ as the upper bound of $\hat{P}$ adaptively. This helps focus inference within a relevant parameter space, avoiding invalid regions to be more efficient. Figure 8 shows this constraint as “scale_p” which is the adaptive range of $\hat{P}$ and “upper_p” is the minimum value of $\hat{P}_{RNA}$. For simplicity, in figures later, we use the original parameter expression instead of the hat expression and it is noted in figure captions as well.
@@ 4.2 Learning result from experimental data
Figure 10: This is a matrix of pairwise plots for Gaussian Process (GP) regression models(Contour plots use a blue color map), illustrating the relationships between parameters and their associated discrepancies. What we see is a slice with GP function values with respect to two variables when all other variables are fixed to certain values and then we see what values these two variables had in all possible data points. The red dots are samples. BOLFI+classifer inference with discrepancy defined in section 3.4 and Figure 8 and learning from experimental data.
@@ 5 Discussions and Conclusion
This study introduces an efficient likelihood-free inference framework for parameter estimation in viral infection dynamics, leveraging the Bayesian Optimization for Likelihood-Free Inference (BOLFI) algorithm. By applying BOLFI to influenza A virus experimental data, we demonstrate that complex mechanistic models with high-dimensional parameter spaces can be inferred accurately and efficiently, even in the absence of explicitly used likelihood functions.
@@
Methodologically, we enhance the BOLFI framework with a GP-based classifier to handle failed cases of simulator and introduce a discrepancy design that integrates multiple types of experimental data. These innovations broaden the applicability of BOLFI to real-world biological inference problems involving heterogeneous data and partial model observability.
@@ Acknowledgments
The project was initiated at iTHEMS RIKEN while Yingying Xu was supported by the SPDR program at RIKEN, Japan. The research process was supported by the Finnish Center for Artificial Intelligence (FCAI). The author thanks Ulpu Remes for helping with classifier design and valuable discussions; Enrico Rinaldi for great support and initial discussions; Catherine Beauchemin and Christian Quirouette for explaining the problem, kindly sharing data and the simulator from their earlier work; Catherine Beauchemin for valuable comments; Henri Pesonen for helpful comments; and Jukka Corander for great support and valuable discussions. Computational resources were provided by CSC – IT Center for Science, Finland, via the Puhti supercomputing infrastructure.